Prev Next

SW&IT Meeting Minutes 2019-09-23

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20190909
LR20190909 - Bec, Matthieu 80.0
LR20190909 - Holmberg, Carl 80.0
LR20190909 - Colson, Andrew 80.0
LR20190909 - Edmeades, David 76.0
LR20190909 - Sola, Igor 40.0
LR20190909 - Kubanek, Petr 95.0
LR20190909 - Hooper, Stephen 40.0
LR20190909 - Zhang, Yang 80.0
LR20190909 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Open topics

  • power interruption Fri 27 (1 to 3 hours)
    • shutdown script readiness ?
  • PEPSI meeting, doodle
  • catalogs.lbto.org
  • iLocater follow-up

Team reports

MB

  • OC, WPC, Management
  • so-nightwriter web/mongodb-ed
  • alta data iterations
  • iLocater, science archive meetings
    • snmp, switch follow-up
  • Pepsi servers discussion
  • catalogs.tucson
next
  • Pepsi system specs
  • Alta smooothed data

DE

  • VMWare support nearly sorted; much internal drama between teams.
  • Colo: rack has arrived, should be installed this week
  • Planning a call with Netops to work on new networks
  • Various user account setups
  • Catalogs.tucson once-over
    • selinux permissive
    • IPTables set to allow only ssh and http
      • Would we have any reason to require https?
    • NFS mount rw; I don't think it needs to be writable
    • Publicly routable IP
    • Is the CGI robust enough for public-facing?
  • Networked KVM switch tests. I think it arrived this morning
  • Temporary storage area on mountain Synology for AO log data
  • Update on date of mountain power switch work?
  • Haven't heard anything from John re: NVidia drivers/Gnome/Xorg. I assume that is good news
  • IDL 7.1 "packaged". It's just a tar of the IDL directory that can be dropped into /usr/local at which point it Just Works. 64-bit binaries.

AC

  • OAC telemetry - added AIPTelemetry.cpp which defines telemetry stream and store values
  • Committed and pushed qwt6 FacSum changes
  • Safety training
next
  • OAC telemetry - create v5.3 branch, test telemetry
  • Create cron job to push and commit and changes to oacontrol.conf
  • Investigate wonky git behavior on oac (v1.7.1)
  • FacSum - remove compass? Pending JH comments.
  • ARGOS planning meeting
  • Safety training

YZ

  • Received feedback from Laura regarding the AO-Cravat POI.
    • Editing the POI to address several of her comments.
    • Remove the design details from the POI into a separate design document.
  • MODS/LBC telemetry:
    • Fixed MODS telemetry ingestion and rebuilt the entire MODS telemetry database, seems to be working fine now.
    • Discussed with Olga about these telemetry. We may need to put more into the database.
    • Working on the LBC telemetry ingestion.
  • MODS/LBC issues:
    • IT#7847, LBC guiding failures for non-sidereal observations. Spent long time on this issue (including a session with Olga). We have some conclusions and discoveries, but not all causes are well understood. Will spend some more time on this issue.
    • IT#7764, MODS "updatepointing" command issue. Reassigned to Rick Poggie. Rick emailed me asking for references and I emailed him back the two current ICE/C++ IIF interface documents.
    • IT#7855, MODS GUI only updates partial target region. This maybe a QT issue. Spent some time looking at the MODS UI code but didn't find anything unusual (but then I do not know much about QT). May next action to Rick to investigate.
    • IT#7854 and #7863, various LBC UI issues, does not seem to happen frequently, may spent some time to look into that.
    • Will schedule some time later with Olga to investigate the LBC FPIA/WRS performance.

PK

  • lots of IT5140 improvements, finished, prepared pull requests
  • finally revealed how to replace following awful, CPU wasting code - patterns page

XZ

Last week:
  • FF matrix quality analyze and document
  • Re-write the actuator analyze idl code
  • Work on the temporart solution for the dump files on ASM comuputer
This week:
  • Analyze the cause of the dump file?
  • Work on the data storage structure
  • Prepare and have a meeting with Arcetri and LBTI about the version control for source code and configration files

IS

Last week:
  • OT
    • Bug fixing and release of OT
  • Start preparing for hotwire talk

This week:

* Prepare Hotwire talk

  • OT
    • LBC calibration header
    • LUCI ARC off lamps
    • Document describing Calibration usage in the XSLT
    • Proposal Table
        • clean proposal table, possibly find out the ones that were approved?
        • add title to table
        • create a list of proposals and emails
  • LUCI
    • add to git /usr/local/binluci_geirs# where # is 1 or 2.
    • look at RTD MOS tab error calculation

CH

Last week:
  • DMS
    • Deployed new scripts and data descriptions enabling individual plots for all Monnit sensors.
    • Loaded Mongodb with all sensor data stored by iMonnit going back to June, 2019.
  • CHAOS
    • Reviewed alert script for MODS1 & 2, isolating temperature and pressure data collection using EPICS and ISIS clients.
  • LBTO/supportscripts dircache
    • Scanned code for use of 'dircache' module removed from Python 3, reviewed alternative directory filename scan functions.

This week:
  • CHAOS
    • Finish reviewing methodology of other non-EPICS alert scripts.
    • Sketch initial concepts for unified alert system.
  • LBTO/supportscripts dircache
    • Discuss exactly what John Hill wants from LBTO/supportscripts/RB_Science.py optimization.
Topic revision: r14 - 24 Sep 2019, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback